Circular No. 8232

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003jb, 2003jc, 2003jd, AND 2003je Four apparent supernovae have been discovered on unfiltered CCD

images: 2003jb by R. Arbour (cf. IAUC 8205), and the rest by LOSS/KAIT (cf. IAUC 8231, via J. Burket, B. Swift and W. Li).

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003jb Oct. 23.84 14^{h}49^{m}22.^{s}28 +63^{o}16′06′′.2 16.5 4′′.6 E, 8′′.0 S
2003jc Oct. 24.2 23 04 53.34 − 6 32 12.1 17.0 20′′.2 W, 6′′.6 N
2003jd Oct. 25.2 23 21 03.38 − 4 53 45.5 16.1 8′′.3 E, 7′′.7 S
2003je Oct. 25.5  8 49 22.24 +36 42 56.6 18.1 3′′.7 W, 19′′.8 N

The above astrometry for SN 2003jb in IC 1065 is by D. Briggs, Portsmouth, England, from an image obtained with a 0.60-m reflector on Oct. 24.911 UT (estimated mag ≈ 15.9); Arbour’s own position end figures, obtained from an image taken in more inferior conditions, are 21^{s}.98, 08′′.6. Arbour adds that SN 2003jb is not present on his image taken on Oct. 17.88 or on Digitized Sky Survey images (limiting magnitudes fainter than 20.0 from a blue plate and 18.5 from a red plate). Additional approximate KAIT magnitudes: SN 2003jc in MCG –01-58-18, Oct. 5.2, [18.5; 25.2, 16.9. SN 2003jd in MCG –01-59-21, Oct. 16.2, [19.0; 26.2, 16.0. SN 2003je in NGC 2668, Oct. 15.5, [19.0; 26.5, 17.9.

S/2003 (22899) 1 W. J. Merline, Southwest Research Institute (SwRI); P. M. Tamblyn,

Binary Astronomy and SwRI; C. R. Chapman, D. Nesvorny, and D. D. Durda, SwRI; C. Dumas, JPL; A. D. Storrs, Towson University; L. M. Close, University of Arizona; and F. Menard, Observatoire de Grenoble, report the discovery on July 26.6 UT, on six direct images (two sets of three images taken 20 min apart in time) made with the Hubble Space Telescope (+ ACS/HRC) in the F606W (600-nm broadband) filter, of a satellite of minor planet (22899) 1999 TO_{14} (V ∼ 18). The satellite is clearly separated in five of these images but streaked in a sixth due to pointing jitter. Trails of several background stars in successive images indicate that the target object is not a background binary star. On July 26.6545, the satellite was at separation 0′′.14 (projected separation 170 km) in p.a. 235^{o}. Using the average albedo of the Koronis family (∼ 0.21), to which (22899) belongs, the size of the primary is estimated to be 4.5 km. The brightness difference is ∼ 2.5 mag, giving an estimated diameter of the secondary of ∼ 1.5 km. This then is the smallest main-belt asteroid known to be binary.

2003 October 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green